What is the length of the altitude of an equilateral triangle whose side has a length of 8?
Answer by richwmiller(17219)

To form the altitude we form a right angle and bisect the base.
so one leg is 1/2 the base which is 1/2*8 the hypotenuse is 8
so we have a^2+b^2=c^2
where a=8/2=4 and c=8
4^2=16
and 8^2=64
so we have 16+b^2=64
solve for b